OpenStudy (anonymous):

Brandon uses 2over3 of a jar of peanut butter each week. He wants to know how many weeks it will take to use 4 jars. Use the model to find 4 ÷ 2over3, the number of weeks it will take.

OpenStudy (anonymous):

dividing by a fraction is the same as multiplying by its reciprocal

OpenStudy (anonymous):

i just need the answer plz then you could get your medal

OpenStudy (anonymous):

ex. \[4\div \frac{3}{7} = 4 \times \frac{7}{3}\]

OpenStudy (anonymous):

i suggest you read the code of conduct we're here to teach, not to give the answer

OpenStudy (anonymous):

if you see anyone just giving out the answers, they're not suppose to and probably do not know any better

OpenStudy (anonymous):

i need to go to bed soon and i need to finish this i have been sick all week and i am better now and this is due tomorrow

OpenStudy (anonymous):

and no one really cares about medals its a lousy incentive

OpenStudy (anonymous):

just read my first and 2nd posts and you should be able to solve it easily

OpenStudy (abb0t):

Answer.

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

who wants the medal answer the thing ...

OpenStudy (abb0t):

Me! Me!!

OpenStudy (anonymous):

give the answer..

OpenStudy (abb0t):

I can't :( http://openstudy.com/code-of-conduct

OpenStudy (dan815):